# Green Edge AI: A Contemporary Survey

### v1

@@ V-A On-device Inference
ICNN  [200] Model selection Image classification ResNet18 (CIFAR-100) MobileNetV2 (CIFAR-100) -$1$3.95% N.A. N.A. 69% 66%
@@
Amortized inference [203] Early termination Image classification ResNet32 (CIFAR-100) ResNet110 (CIFAR-100) +3.8% -0.87% $-75.05\%$${}^{\text{b}}$ $51.02\%$ N.A. N.A.
BranchyNet [204] Early termination Image classification AlexNet (CIFAR-10) ResNet110 (CIFAR-10) +0.81% -1.53% $33.9\%\ (58.7\%)$${}^{\text{c}}$ $46.4\%\ (47.5\%)$ N.A. N.A.
FlexDNN [206] Early termination Activity recognition Inception-V3 (UCF-15) VGG-16 (UCF-15) $+3\%$ $+7\%$ $3.9$×$\times\ (69.0\%)^{\text{d}}$ $2$×$\times\ (38.1\%)$ $73.1\%$ $50.9\%$
@@
CGNet [209] Layer/channel skipping Image classification ResNet18 (CIFAR-10) VGG16 (CIFAR-10) -0.56$\sim$-0.04-0.39$\sim$+0.085.475.475.475.47$\times$×$\sim$7.957.957.957.95$\times$${}^{\text{b}}$ $3.41$$\times$×$\sim$5.105.105.105.10$\times$ N.A. N.A.
@@
  • a: The accuracy is measured by the peak signal-to-noise ratio (PSNR), and the computation saving is measured by the speedup when MobiSR is implemented on the Qualcomm SDM845 platform [201].
  • b: The computation saving is measured by FLOP reduction.
@@
  • c: The computation saving is measured by the speedup on a platform with a 3.0 GHz CPU (NVIDIA GeForce GTX TITAN X (Maxwell) GPU).
@@
  • d: The computation saving is measured by the speedup (frame drop rate reduction) on a Samsung Galaxy S8 smartphone.
V-B MEC Server-based Inference
@@
1) Commmunication-Computational Resource Management: Given the inference data to be offloaded, green communication techniques can be adopted to design energy-efficient uplink transmission schemes [212]. For EUDs that are able to compute some inference tasks, inference task offloading policies should be jointly designed. Meanwhile, efficient computational resource management at the MEC server is crucial to accomodate more inference requests to achieve maximum on-device energy savings. Notably, these considerations align with those in conventional MEC systems [10].
@@
Nevertheless, task offloading in MEC systems mainly aims at reducing the execution latency and energy consumption, while inference accuracy becomes a new performance metric in MEC server-based inference [213]. To narrow the research gap, an inference task distribution scheme was presented in [214], which considers both inference accuracy and computation demand of DNN models using statistical information estimated on a validation dataset. In addition to computation overhead, the placement overhead of DNN models at the MEC server is also non-negligible. To maximize the average inference accuracy while meeting the latency requirements, a multi-dimensional optimization problem was formulated and efficiently solved by DRL in [215], which determines the DNN placement and communication-computational resource allocation. A recent investigation in [216] further considers the purchase of carbon emission rights when designing inference task offloading policies.
@@
2) Data Preprocessing: Many DNN models accept input data with different configurations. For example, the real-time object detector YOLO [217] can process video frames in different resolutions. Some data configurations, such as frame rate, resolution, and encoding bitrate, are tightly coupled with the communication overhead [218]. Thus, preprocessing the raw inference data to proper configurations has great potential for improving the EE of MEC server-based inference [219].
@@
The impact of data preprocessing on inference accuracy can be estimated via offline profiling. For a bearing vibration signal dataset, the facility fault diagnosis accuracy of AlexNet was found increasing with the sampling rate [220]. Since a higher sampling rate leads to an increased data volume, it needs to be dynamically configured according to the wireless channel condition. Lossy compression was exploited to reduce the communication overhead on an MNIST classification task in [221], where the inference accuracy was modeled as a function of compression ratio via a lookup table. In [218], the frame quality for video analytics, including frame rate, resolution, and bitrate, was controlled considering both the energy cost and inference accuracy under a data-driven framework. Similarly, joint configuration adaptation and bandwidth allocation for multi-user video analytics systems was investigated in [222], where the inference accuracy was modelled as a concave function with respect to resolution and frame rate. The data configuration in MEC server-based inference can also be adapted without finding an explicit function of the inference accuracy, which was achieved via DRL and Bayesian online learning in [223] and [224], respectively. In [225], the DNN placement, inference task offloading, data preprocessing, and communication-computational resource allocation are jointly optimized for industrial IoT applications.
@@ V-C Device-edge Co-inference
For on-device inference, DNN models are fully executed at EUDs that may drain battery energy quickly. For MEC server-based inference, uploading the possibly high-dimensional data is both time- and energy-consuming. To maximize the benefits of both paradigms, device-edge co-inference, which partitions a backbone DNN model between an EUD and the MEC server [14], appears to be a more favorable solution.
@@
As shown in Fig. 5, the on-device DNN partition (i.e., the first few layers of the backbone DNN) extracts a feature map from the input for transmission to the MEC server. The MEC server executes the server DNN partition (i.e., the remaining layers of the backbone DNN) using the received feature map to derive the final inference result. Nevertheless, realizing energy-efficient device-edge co-inference is beyond model partitioning and requires additional optimization as discussed in the sequel.
@@
1) Adaptive Model Partitioning: A study in [226] measures the computation and data overhead of various DNN models at all candidate model partition points. Based on such information, a lightweight scheduler named Neurosurgeon was developed to select the best partition point in runtime. Compared to MEC server-based inference, Neurosurgeon reduces $59.5\%$ of the mobile energy consumption. Also, the backbone DNN in device-edge co-inference can be a branchy network [227, 228, 229]. In particular, Edgent [227] exploits both model partitioning and early termination, where the optimal model partition point and early exit are determined via a configuration map for dynamic network status. While the early exit of Edgent is fixed for all input once selected, SPINN [228] terminates the inference process at different exits in adaptation to the input data. Nevertheless, the large search space of candidate model partition points and early-exit confidence score thresholds necessitates low-complexity optimization. Therefore, EdgeML [229] trains a light DRL agent that determines the optimal configurations of a branchy network according to system dynamics, which takes accuracy, latency, and energy consumption into account.
@@
2) Feaure Compression: Model partitioning amortizes the computation workload between EUDs and the MEC server. However, the communication overhead may be increased for certain DNN models at some partition points because of the in-layer data amplification phenomenon [226, 230]. To achieve energy-efficient device-edge co-inference, it is thus necessary to reduce the feature size by exploiting the statistical characteristics [231]. Empirical results show lossless compression can achieve a $1.5$$\sim$3×3$\times$3 × compression ratio, while lossy compression eliminates the redundancy more aggressively with some accuracy loss [232]. In [233], feature compression in device-edge co-inference was performed by an autoencoder, which is adapted to the channel condition to strength the robustness. Besides, the EE-oriented design methods for on-device and MEC server-based inference, e.g, model compression [51, 234, 235, 236], early termination [237, 238], and adaptive communication [239, 240], can be jointly utilized with feature compression. In addition, a theoretical characterization of the communication overhead-inference accuracy tradeoff is also critical. This was achieved in [241] via the information bottleneck principle [242], and based on which, a task-oriented communication scheme that only transmits suffcient but minimal information relevant to the inference task was developed.
@@
When EUDs make independent and concurrent inference requests, model parititioning and computational resource management should be jointly optimized [248, 249, 250, 251, 252]. For example, with a data-driven multi-core CPU model, the optimal model partitioning and computational resource allocation plan was obtained via an alternating optimization algorithm in [248]. For applications where the inference requests arrive sequentially, an online inference request admission policy was developed based on reinforcement learning in [249] to minimize the energy consumption of both EUDs and MEC servers. In [250], multiple efficient inference techniques, including batching and early termination, were optimized jointly with the allocation of commmunication and computational resources. Decentralized multi-agent reinforcement learning-based energy-efficient collaborative inference schemes were developed in [251], which exploits the learning experience of neighoring EUDs to accelerate policy optimization in large-scale edge AI systems. A prototype showed that more than $40\%$ of the mobile energy consumption and inference latency can be saved compared with a baseline scheme. Assuming an autoencoder-based feature compression scheme, the model partitioning, and channel and power allocation of multiple EUDs were jointly optimized in [252], which saves up to $56\%$ of inference latency and $72\%$ of energy consumption.
@@ V-D Takeaways
This section reviews the EE-oriented design approaches for three typical edge inference paradigms, including on-device inference, MEC server-based inference, and device-edge co-inference. For on-device inference, the essence is to reduce the computational energy consumption by removing the less important processing of DNNs using model compression, energy-aware NAS, and adaptive inference techniques. Although efficient processing of DNNs has been investigated for years, there is still a genuine need to modify the existing solutions for diverse edge AI platforms with strict computational resource and energy constraints. For MEC server-based inference, it is essential to develop energy-efficent task offloading strategies, which reduces both the on-device computation energy consumption and transmission energy consumption. These strategies are largely inherited from those developed for conventional MEC systems, yet the distinctive fault-tolerence property of DNN workloads should also be considered via data preprocessing. The device-edge co-inference paradigm provides an agile means to deploy DNN models accross multiple computing nodes in edge AI systems. Similar to MEC server-based inference, removing the redundant communication overhead in intermediate feature maps is critical to save the transmission energy. It is also critical to strike a good balance among the on-device computation workload, communication overheand, and inference accuracy. In addition, the potential collaboration among EUDs can not only harvest the distributed computational resources, but also benefit from the task dependence and spatial data correlation. We foresee in future edge AI systems, the three edge inference paradigms coexist for heterogeneous applications. Therefore, unifying their respective EE-oriented design methodologies would be necessary.
@@ VI Future Research Directions
In this section, we highlight several prospective future research directions to enable green edge AI. Specially, we first introduce the low-power sensing techniques via integrated sensing and communication (ISAC). Hardware/software co-designs for edge AI are then discussed, which complement the existing algorithmic innovations to further boost the EE. Besides, we elaborate two critical aspects of neuromorphic computing for ultra-energy-efficient edge AI, including the spiking neural networks (SNNs) and compute-in-memory (CIM) techniques. In addition, green energy-powered edge AI that can potentially achieve carbon-neutral operations, is also investigated.
@@ VI-A Low-Power Sensing via ISAC
With the seamless integration of sensing and communication, 6G wireless networks are poised to become perceptive, essentially opening their “eyes” [253] for numerous intelligent applications, including high-accuracy localization and tracking, simultaneous localization, imaging, and mapping, augmented human sensing, as well as gesture and activity recognition. A key success of ISAC lies in empowering edge AI systems to “see” the physical world by exploiting radio wave propagation. Particularly, MEC servers can leverage communication signals for device-free sensing [254], liberating EUDs from the power-intensive data acquisition tasks in centralized edge learning. To support ISAC, full-duplex MEC servers are commonly deployed. However, this setup poses new challenges, as both the echo sensing signals from sensing targets and the uplink transmission signals from EUDs may arrive before the downlink transmission concludes. The severe self-interference necessitates a delicate scheme to manage mutual interference between sensing and communication signals [255].
On the other hand, EUDs can also reuse the uplink/sidelink signals for low-power sensing, through either multiple access or D2D channels [256]. In each ISAC frame, EUDs transmit the sensed data from the previous frame while simultaneously sense the environment. Nevertheless, a challenge arises as the conventional uplink transmission beams are usually directed toward the receiver (i.e, the MEC server in edge AI systems), which are not optimal for surrounding sensing. Consequently, an effective waveform and beamforming design is essential [257]. Another crucial aspect in ISAC-enabled low-power sensing is the trade-off between sensing accuracy, communication rate, and centralized edge learning performance [258], which needs proper understanding and management to achieve the desired outcomes in green edge AI systems.

### v2

@@ V-B MEC Server-based Inference
ICNN [220] Model selection Image classification ResNet-50 (CIFAR-100) MobileNetV2 (CIFAR-100) $-11.11\%$ $-13.95$% $22\%$ ($64\%$)${}^{\text{a}}$ N.A. $69\%$${}^{\text{a}}$ $66\%$
@@
Cascaded inference [221] Early termination Image classification ResNet-32 (CIFAR-100) ResNet-110 (CIFAR-100) $+3.8\%$ $+0.87\%$ $-75.05\%$${}^{\text{c}}$ $51.02\%$ N.A. N.A.
BranchyNet [222] Early termination Image classification AlexNet (CIFAR-10) ResNet-110 (CIFAR-10) $+0.81\%$ $-1.53\%$ $33.9\%\ (58.7\%)$${}^{\text{d}}$ $46.4\%\ (47.5\%)$ N.A. N.A.
@@
[…]
  • a: The computation saving is measured by the speedup on a CPU (Eyeriss) platform. When with restrictions on the accuracy loss, the amount of energy saving generally reduces [220].
@@
  • b: The accuracy is measured by the peak signal-to-noise ratio (PSNR), and the computation saving is measured by the speedup when MobiSR is implemented on the Qualcomm SDM845 platform [227].
  • c: The computation saving is measured by FLOP reduction.
@@
  • d: The computation saving is measured by the speedup on a platform with a 3.0 GHz CPU (NVIDIA GeForce GTX TITAN X (Maxwell) GPU) [222].
@@
  • e: The computation saving is measured by the speedup (frame drop rate reduction) on a Samsung Galaxy S8 smartphone [229]. The energy saving is calculated for processing each frame.
  • f: The energy saving is measured on a Digilent ZedBoard Zynq-7000 development board [232].
@@
1) Communication-Computational Resource Management: Given the inference data to be offloaded, green communication techniques can be adopted to design energy-efficient uplink transmission schemes [67, 234]. For EUDs that are able to compute some inference tasks, inference task offloading policies should be jointly designed. Meanwhile, efficient computational resource management at the MEC server is crucial to accommodate more inference requests to achieve maximum on-device energy savings. Notably, these considerations align with those in conventional MEC systems [10].
@@
Nevertheless, task offloading in MEC systems mainly aims at reducing the execution latency and energy consumption, while inference accuracy becomes a new performance metric in MEC server-based inference [34, 235]. To narrow the research gap, an inference task distribution scheme among EUDs and the MEC server was presented in [236], which considers both inference accuracy regarding data quality and computation demand of DNN models using statistical information estimated on a validation dataset. In addition to computation overhead, the placement overhead of caching DNN models at the MEC server is also non-negligible. To maximize the average inference accuracy while meeting the latency requirements, a multi-dimensional optimization problem was formulated and efficiently solved by DRL in [237], which determines the DNN placement and communication-computational resource allocation. A recent investigation in [238] further considers the purchase of carbon emission rights when designing inference task offloading policies.
@@
2) Data Preprocessing: Many DNN models accept input data with different configurations. For example, the real-time object detector YOLO [239, 240] can process video frames in different resolutions. Some data configurations, such as frame rate, resolution, and encoding bitrate, are tightly coupled with the communication overhead [241]. Thus, preprocessing the raw inference data to proper configurations has great potential for improving the EE of MEC server-based inference [242].
@@
The impact of data preprocessing on inference accuracy can be estimated via offline profiling. For a bearing vibration signal dataset, the facility fault diagnosis accuracy of AlexNet was found to increase with the sampling rate [243]. Since a higher sampling rate leads to an increased data volume, it needs to be dynamically configured according to the wireless channel condition. Lossy compression was exploited to reduce the communication overhead on the MNIST and CIFAR-10 classification tasks in [244], where the inference accuracy was modeled as a function of compression ratio via a lookup table. In [241], the frame quality for video analytics, including frame rate, resolution, and bit rate, was controlled considering both the energy cost and inference accuracy under a data-driven framework. Similarly, joint configuration adaptation and bandwidth allocation for multi-user video analytics systems was investigated in [245], where the inference accuracy was modeled as a bi-concave function with respect to resolution and frame rate. The data configuration in MEC server-based inference can also be adapted without finding an explicit function of the inference accuracy, which was achieved via DRL and Bayesian online learning in [246] and [247], respectively. In [248], the DNN placement, inference task offloading, data preprocessing, and communication-computational resource allocation are jointly optimized for industrial IoT applications.
@@ V-C Device-edge Co-inference
For on-device inference, DNN models are fully executed at EUDs that may drain battery energy quickly. For MEC server-based inference, uploading the possibly high-dimensional data is both time- and energy-consuming. To maximize the benefits of both paradigms, device-edge co-inference, which partitions a backbone DNN model between an EUD and the MEC server [14], appears to be a more favorable solution. It can be envisaged that device-edge co-inference can improve both time and energy efficiency by harnessing computational resources at different nodes, despite extra communication and computation overhead compared to on-device and MEC server-based inference, respectively.
@@
As shown in Fig. 6, the on-device DNN partition (i.e., the first few layers of the backbone DNN) extracts a feature map from the input for transmission to the MEC server. The MEC server executes the server DNN partition (i.e., the remaining layers of the backbone DNN) using the received feature map to derive the final inference result. In the device-edge co-inference paradigm, there exists a general trade-off between the communication and computation energy consumption for a target inference accuracy, which can be adjusted via model partition point selection [56]. Nevertheless, realizing energy-efficient device-edge co-inference is beyond model partitioning and requires additional optimization as discussed in the sequel.
@@
1) Adaptive Model Partitioning: A study in [249] measures the computation and data overhead of various DNN models at all candidate model partition points. Based on such information, a lightweight scheduler named Neurosurgeon was developed to select the best partition point in runtime. Compared to MEC server-based inference, Neurosurgeon reduces $59.5\%$ of the energy consumption on an EUD. The backbone DNN model in device-edge co-inference can also be a branchy network [250, 251, 252]. In particular, Edgent [250] exploits both model partitioning and early termination, where the optimal model partition point and early exit are determined via a configuration map for dynamic network status. While the early exit of Edgent is fixed for all input once selected, SPINN [251] terminates the inference process at different exits in adaptation to the input data. Nevertheless, the large search space of candidate model partition points and early-exit confidence score thresholds necessitates low-complexity optimization. Therefore, EdgeML [252] trains a light DRL agent that determines the optimal configurations of a branchy network according to system dynamics, which takes accuracy, latency, and energy consumption into account.
@@
2) Feature Compression: Model partitioning amortizes the computation workload between EUDs and the MEC server. However, the communication overhead may be increased for certain DNN models at some partition points because of the in-layer data amplification phenomenon [249, 253]. To achieve energy-efficient device-edge co-inference, it is thus necessary to reduce the feature size by exploiting statistical characteristics [254]. Empirical results show that lossless compression can achieve a $1.5$$\sim$3×3$\times$3 × compression ratio, while lossy compression eliminates the redundancy more aggressively with some accuracy loss [255]. In [256], feature compression in device-edge co-inference was performed by an autoencoder, which is adapted to the channel condition to strength the robustness. Besides, the EE-oriented design methods for on-device and MEC server-based inference, e.g, model compression [56, 257, 258, 259], early termination [260, 261], and adaptive communication [262, 263], can be jointly utilized with feature compression. In addition, a theoretical characterization of the communication overhead-inference accuracy trade-off is also critical. This was achieved in [264] via the information bottleneck principle [265], and based on which, a task-oriented communication scheme that only transmits sufficient but minimal information relevant to the inference task was developed.
@@
When EUDs make independent and concurrent inference requests, model partitioning and computational resource management should be jointly optimized [273, 274, 275, 276, 277]. For example, with a data-driven multi-core CPU model, the optimal model partitioning and computational resource allocation plan was obtained via an alternating optimization algorithm in [273]. For applications where the inference requests arrive sequentially at EUDs, an online inference request admission policy was developed based on reinforcement learning in [274] to minimize the energy consumption of both EUDs and MEC servers. In [275], multiple efficient inference techniques, including batching and early termination, were optimized jointly with the allocation of communication and computational resources. Decentralized multi-agent reinforcement learning-based energy-efficient collaborative inference schemes were developed in [276], which exploits the learning experience of neighboring EUDs to accelerate policy optimization in large-scale edge AI systems. A prototype with two EUDs and two MEC servers showed that more than $40\%$ of the mobile energy consumption and inference latency can be saved compared with a baseline scheme. Assuming an autoencoder-based feature compression scheme, the model partitioning, and channel and power allocation of multiple EUDs were jointly optimized in [277], which saves up to $56\%$ of inference latency and $72\%$ of energy consumption.
@@ V-D Takeaways
This section reviews the EE-oriented design approaches for three typical edge inference paradigms, including on-device inference, MEC server-based inference, and device-edge co-inference. For on-device inference, the essence is to reduce the computation energy consumption by removing the less important processing of DNNs using model compression, energy-aware NAS, and input adaptive inference techniques. Although efficient processing of DNNs has been investigated for years (e.g., [8] and references therein), there is still a genuine need to modify the existing solutions for diverse edge AI platforms with strict computational resource and energy constraints. For MEC server-based inference, it is vital to develop energy-efficient inference task offloading strategies, which reduces both the on-device computation energy consumption and transmission energy consumption. These strategies are largely inherited from those developed for conventional MEC systems, yet the distinctive fault-tolerance property of DNN workloads should also be considered via data preprocessing. The device-edge co-inference paradigm provides an agile means to deploy DNN models across multiple computing nodes in edge AI systems. Similar to MEC server-based inference, removing the redundant communication overhead in intermediate feature maps is critical to save the transmission energy. It is also mandatory to strike a good balance among the on-device computation workload, communication overhead, and inference accuracy. In addition, the potential collaboration among EUDs can not only harvest the distributed computational resources, but also benefit from the task dependency and spatial data correlation. We foresee in future edge AI systems, the three edge inference paradigms coexist for heterogeneous applications. Therefore, unifying their respective EE-oriented design methodologies and unraveling scaling laws of energy consumption regarding the EUD and MEC infrastructure densities would be necessary. Moreover, in applications where sampling the inference data incurs significant overhead [75], understanding the impact of sensing on edge inference becomes crucial.
@@ VI Future Research Directions
In this section, we highlight several prospective future research directions of green edge AI. Specially, we first introduce the low-power sensing techniques via integrated sensing and communication (ISAC). Hardware/software co-designs for edge AI are then discussed, which complement the existing algorithmic innovations to further boost the EE. Besides, we elaborate two critical aspects of neuromorphic computing for ultra-energy-efficient edge AI, including the spiking neural networks (SNNs) and compute-in-memory (CIM) techniques. In addition, green energy-powered edge AI that can potentially achieve carbon-neutral operations, is also investigated. Finally, we elaborate several possible approaches for delivering generative AI (GenAI) services at mobile edge networks in an energy-efficient manner.
@@ VI-A Low-Power Sensing via ISAC
With the seamless integration of sensing and communication, 6G wireless networks are on the verge of revolutionizing edge AI systems, enabling them to perceive and understand the physical world in unprecedented ways [278]. These advancements can benefit a plethora of emerging applications, including high-accuracy tracking, simultaneous localization, imaging, and mapping, augmented human sensing, as well as autonomous vehicles and the Metaverse.
The key success of ISAC lies in empowering edge AI systems to “see” the environment by exploiting radio wave propagation. In this context, MEC servers play a prominent role by utilizing communication signals for device-free sensing [279]. By doing so, EUDs are liberated from the power-intensive data acquisition tasks that would otherwise burden the centralized edge learning process. To support ISAC, full-duplex MEC servers need to be deployed. However, this setup introduces new challenges due to the simultaneous arrival of echo sensing signals from sensing targets and uplink transmission signals from EUDs before the downlink transmission concludes. The resulting severe self-interference necessitates the developments of dedicated schemes to effectively manage mutual interference between sensing and communication signals [280].
On the other hand, EUDs in edge AI systems can leverage the uplink/sidelink signals for low-power sensing, utilizing either multiple access or D2D channels [281]. The sensed data can be offloaded to an MEC server for downstream inference following the novel integrated sensing, computation, and communication (ISCC) paradigm [282]. In each ISAC frame, EUDs transmit the sensed data from the previous frame while simultaneously sensing the environment. Nevertheless, a critical issue arises as the conventional uplink transmission beams are typically directed toward the receiver (i.e., the MEC server in edge AI systems), which may not be optimal for sensing the surroundings. Therefore, effective waveform and beamforming designs specifically tailored for edge AI applications are essential to ensure accurate sensing and efficient data offloading [283]. Another important aspect in ISAC-enabled low-power sensing for edge AI is the trade-off between sensing accuracy, communication rate, and intelligence performance [284]. Proper understanding and management of this triple trade-off are necessary. By optimizing the related parameters, ISAC can empower a broad range of applications with energy-efficient and accurate perception capability to achieve green edge AI.